IODP Proceedings    Volume contents     Search



After Hole U1373A, on the eastern side of the summit plain of Rigil Guyot, had to be abandoned, the vessel was offset in dynamic positioning mode to Site U1374, on the western side of the summit plain. The 5.6 nmi offset was accomplished in 3.25 h, and by 1630 h on 5 January 2011 the ship was positioning at the new location (all times are New Zealand Daylight Time, Universal Time Coordinated [UTC] + 13 h). The vibration-isolated television (VIT) frame was deployed and used to monitor the bit tagging the seafloor at 1570.0 meters below rig floor (mbrf) (1559.0 meters below sea level [mbsl]), 2.6 m shallower than the corrected precision depth recorder depth.

Hole U1374A was spudded with a Type C-4 rotary core barrel bit at 2035 h on 5 January (Fig. F3; Table T1). After penetrating a thin sedimentary cover, the bit penetrated igneous basement at 16.7 mbsf. By 0345 h on 10 January, rotary coring had slowly advanced to 130.4 mbsf, with an average recovery of 84%. At this time, the bit had acquired 81.3 rotating hours, and preparations were made to recover the drill string and replace the bit with a fresh unit. Penetration into basement was 113.7 m, with an average recovery of 85%. The average rate of penetration while coring basement was 1.6 m/h.

A free-fall funnel (FFF) was made up and deployed at 0550 h on 10 January. The VIT frame was launched, and the bit extraction from the hole was monitored with the VIT system to ensure that the FFF was not dislodged during the process. The used bit was at the rotary table at 1145 h and was replaced with a new Type C-4 bit. The latch sleeve was also inspected, and the mechanical bit release (MBR) was replaced with a rebuilt unit. The used bit was found to be in excellent condition and only inch undergauge, with all inserts intact and very minor wear across all rows. The bearings were also still moderately tight.

The FFF was reentered at 1635 h. The bottom-hole assembly (BHA) was run in the hole without incident while the driller maintained slow rotation and a low pump rate. There was no fill at the bottom of the hole. Coring resumed at 1900 h on 10 January. During the course of the week the scientific party decided to core with the present bit until its destruction and to forego a second bit trip at this site. Rotary coring continued until 0545 h on 18 January, when erratic pump pressure indicated that there was a problem at the bottom of the drill string. Instead of continuing to core with a bit that had far outlasted its expected life at 133.4 rotating hours, it was decided to end coring at a final depth of 522.0 mbsf. Penetration into basement was 505.3 m, with an average recovery and rate of penetration of 88.0% and 2.5 m/h, respectively. The average recovery for the entire hole was 87.8%, with an average penetration rate of 2.4 m/h.

After a routine wiper trip, which suggested that the hole was in good condition, an attempt was made to release the bit with the rotary shifting tool (RST). There was no positive indication that the window sleeve in the MBR was shifted, and it was initially thought that an obstruction in the throat of the bit was preventing the tool from engaging the window latch. To possibly remedy this, a bit deplugger was pumped down, but there was no indication by a change in pump pressure that it had landed. The deplugger had to be jarred free and was recovered on the coring line. Multiple abrasions on the deplugger indicated that it was mechanically prevented from landing in the throat of the bit. Because downhole logging in this particular deep hole was considered important, it was decided to trip the drill pipe, remove the bit and MBR, and make up a shorter logging BHA with a 9¼ inch diameter logging/clean-out bit that would allow more of the open hole to be logged.

As the drill pipe was recovered, sepiolite-laden mud was discharged under pressure on the rig floor with each connection. In order to remedy this, the circulating head was made up and the pipe was flushed out. It is possible that during this process the bit may have been released and fallen to the seafloor, because no further mud was discharged as the remaining pipe was recovered. The exact time and position of the bit release can only be speculated.

The end of pipe cleared the rotary table at 2015 h on 18 January. A shorter logging BHA was made up with the cleanout bit and deployed at 0145 h. The bit entered the FFF at 0430 h and was positioned at a logging depth of 101.2 mbsf. The standard triple combination (triple combo) suite was made up and run in at 1045 h. The tool was not able to advance into the open hole because of a bridge ~7 m below the end of the drill pipe. The tool was recovered, and the bit was lowered to 143.6 mbsf to clear the bridge. The bit was then pulled back and placed at 110.8 mbsf. When the triple combo tool string was unable to advance below 138.0 mbsf, the tool string was again recovered and reconfigured to a shorter assembly using only the density and gamma ray sensors in the hope that this tool string would more easily negotiate hole entry. After this attempt was unsuccessful, the drill pipe was recovered with the bit, clearing the rotary table at 1115 h.

A four-stand rotary core barrel coring assembly was made up with a new MBR and a used Type C-4 bit (recovered from Hole U1373A). The drill string entered the FFF at 1540 h on 20 January. From 1730 to 2315 h the hole was washed and reamed to bottom (522 mbsf), flushed with mud, and then displaced with 165 bbl of heavy (10.5 ppg) mud. The bit was dropped at the bottom of the hole, and the end of pipe was placed at a logging depth of 128.1 mbsf, ~18 m deeper than the previous logging attempt and below a potentially unstable zone in the formation. An extra stand of drill collars was added to the BHA to keep the tapered drill collar as close to the seafloor as possible, as that is the usual choke point at which a BHA becomes stuck.

The triple combo logging tool string was made up (in its standard extended version) and deployed again at 0540 h on 21 January and succeeded in logging the hole up from 520 mbsf. The tool suite was recovered and laid out. The second instrument deployed was the Göttingen Borehole Magnetometer (GBM), which made one full pass down from the rig floor to 520 mbsf and back up. Communication with the GBM was lost while it was being retrieved in the pipe; however, because a sighting of the tool was carried out at the start of deployment, the rotation history of the GBM may still be obtainable. The third logging run was performed with the Formation MicroScanner-sonic tool string, which also successfully came within 2 m of the bottom of the hole. The fourth log was conducted with the Ultrasonic Borehole Imager. The fifth log was a redeployment of the GBM tool. All runs were successful.

While the GBM was being deployed for the second time, the driller noticed that the suspended string weight was decreasing slightly, indicating that the formation was starting to squeeze the BHA. In order to compensate for this, the driller picked up the string an additional 3 m. While the logging tool was being retrieved the vessel had to be offset 75 m in order to lower the pipe connection the additional 3 m so that the slips could be set. From 2215 h on 22 January to 0930 h on 23 January the driller attempted to free the drill string. Although circulation was maintained, there was no rotation and the drill pipe could not be raised or lowered. Realizing that further efforts would be fruitless, the crew made the necessary preparations to sever the drill string directly above the tapered drill collar. The top of the tapered drill collar was at ~13 mbsf. The drill pipe was severed at 1500 h on 23 January.

Once the drill pipe was recovered and the beacon was retrieved, the vessel was secured for sea. The vessel departed for Site U1375 at 1945 h on 23 January, with an estimated time of arrival of 0500 h on 25 January.